OpenStudy (anonymous):

Suppose a magnetic field of 0.20 T is applied to the area within a circular loop of wire. The strength of the field is then changed to 0.10 T. By what factor is the magnetic flux changed? 1/2 , 1/4 , 2 , or 4 ? :/

OpenStudy (michele_laino):

before variation, the magnetic flux is: 0.20 * S, where S is the area enclose by the circular loop after the variation, the magnetic flux is: 0.10*S so the magnetic flkux changing is: \[\Delta \Phi = 0.2S - 0.1S = 0.1S\]

rvc (rvc):

do you know the relation between the two?

OpenStudy (michele_laino):

and the requested ratio is: \[\frac{{{\Phi _2}}}{{{\Phi _1}}} = \frac{{0.1S}}{{0.2S}} = ...\]

OpenStudy (anonymous):

ohh okay! so then we get 1/2 as our solution?

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (anonymous):

yay! thanks!!:D

OpenStudy (michele_laino):

:)